# Anatomy of an ionized bubble: NIRCam grism spectroscopy of the double-peaked Lyman- emitter COLA1 and its environment

### v1

@@ 3.2 Redshift and line fluxes measurements
In order to measure the line fluxes and redshifts, we extract the 1D spectra of the candidates. We use the [O III]λ5008 line to trace the morphology of each object, as it is the most luminous emission line in our range. First, we select regions in the grism 2D EMLINE spectra containing the [O III]λ5008 line, and integrate them along the wavelength direction, for both modules A and B, whenever they are available. Then, we fit the spatial profile of the [O III]λ5008 emission with a single Gaussian for most cases, and double or triple Gaussian components for a few objects that present multiple clumps compatible with [O III] emission at the given redshift. Finally, we optimally extract the 2D spectra with the shape derived from the Gaussian fit to obtain the 1D spectrum of each candidate.
@@
Lastly, we measure the line fluxes. We fit individual Gaussian profiles to [O III]λ5008, [O III]λ4960 and H. We use narrow flat priors for the observed wavelengths of each line, based on the redshifts extracted from the [O III] doublets. In the particular case of COLA1, we also measure fluxes of the [O III]λ4364 and H, due to their exceptional S/N, as detailed in Sect. 3.2.1 below.
3.2.1 Redshift calibration
@@ 3.3 SED modeling
We performed spectrophotometric SED fitting of all the [O III] emitters in our sample using the Prospector code (Johnson et al., 2021), including nebular emission and nebular continuum implemented via Cloudy (Ferland et al., 1998; Byler et al., 2017) and FSPS (Conroy et al., 2009, 2010; Conroy & Gunn, 2010) with the MIST stellar models (Choi et al., 2017). We follow the setup described in Naidu et al. (2022b), which in turn is adapted from Tacchella et al. (2022). We fit the F115W, F150W, F200W and F356W photometric fluxes, as well as H and [O III]λ5008,4960 line-fluxes simultaneously. The free parameters in the model include seven bins for the non-parametric star-formation history, with the first two being fixed to 0–5 Myr and 5–10 Myr (e.g., Tacchella et al. 2023), and the remaining logarithmically spaced out to . The other free parameters capture the total stellar mass formed, the stellar metallicity, the gas-phase metallicity, dust attenuation (a screen and an additional component for young stars), and the ionization parameter. We assume a Chabrier (2003) initial mass function.
@@
Figure 5 shows the best-fit spectrum, the predicted photometry, and draws from the posterior that are compared to the observed data. The model predicts a best-fitting UV slope of , which is at odds with the UV photometry that shows a peculiar ”U” shape where the F150W flux lies below F115W and F200W (see Sect. 4.3). Keeping in mind that significant LyC leakage is at play, we fit a model with additional parameters that capture the escape fraction (Conroy & Kratter, 2012). In particular, a fraction of OB stars (uniform prior between 0 to ) is allowed to shine through the birth-cloud and dust-screen entirely unimpeded, which mimics “escape through holes”, i.e., an ionization-bounded nebula (e.g., Fig. 1 in Zackrisson et al. 2013). The predicted escape fraction from this fit is %. This high of is in agreement with the values obtained through different observational estimators in Sect. 4.5 below. However, due to the sparse data, we note the quality of our fit is effectively indistinguishable from a fit assuming and we defer investigation of this issue to future work, in particular when more photometric bands are available. This exploration highlights how COLA1 provides a unique test case, directly in the EoR, to hone estimators of LyC .
@@
COLA1 shows an average EW([O III]+H) when compared with typical high-redshift galaxies in the EIGER sample (Matthee et al., 2023a). Using the grism fluxes of the [O III]+H lines and subtracting them from the F356W photometry to get a direct estimation of the continuum we get a value of EW0([O III]+H) Å, which is very consistent, with the Prospector prediction of EW0([O III]+H) Å. The fitted stellar mass is , but we note that these errors may be underestimated as they are partly bounded by the range in models that are included in Prospector. For example, if we would allow more recent and or exotic models, such as new implementations of binary star physics (Lecroq et al., 2024) or a top heavy initial mass function (e.g., Cameron et al., 2023a), it may well be possible that a somewhat lower stellar mass and star formation rate is required to reproduce COLA1s spectrum.
@@ 3.4 Galaxy mock catalogs in the COLA1 Field
In order to identify and quantify over-densities of the detected [O III] emitters in our data (shown in Fig. 6), we need to account for the selection function. As illustrated in Fig. 1, the field of view and sensitivity are both non-trivially dependent on wavelength. Moreover, another obvious factor in our selection function is that we targeted JWST to center on the bright galaxy COLA1. Because of these complications, we compare our observed distribution of galaxies to detailed mock observed samples of [O III] emitters that are derived by forward-modeling simulations through our selection function (i.e. the error cube that is illustrated in bottom panel of Fig. 1). This methodology has been developed by Mackenzie et al. (in prep.) in the context of understanding the environments of luminous quasars in the EIGER survey (see also Eilers et al. 2024), but is easily applicable to our survey.
As detailed in Mackenzie et al. (in prep.), large N-body simulations are combined with the UniverseMachine model (Behroozi et al., 2019) to obtain the SFR of each simulated halo. Then, by matching the rank order of the SFR and the [O III] luminosity (the property on which we select galaxies), one can derive an empirical description to obtain the [O III] luminosity of each simulated halo that is calibrated to match the observed [O III] luminosity function at . The observed [O III] luminosity function is an updated version of the one presented in Matthee et al. (2023a), with minor changes. Since the clustering strength and the [O III] to UV luminosity relations of the simulated galaxies match the observed counterparts, this method seems to place the right galaxies in the right halos. The method is applied to the Uchuu (Ishiyama et al., 2021) dark matter only simulation. The Uchuu simulation is among the largest, (2 Gpc/)3, and therefore contains the rarest structures and large samples of halos hosting galaxies like COLA1, at the expense of resolution (i.e. it only resolves galaxies with [O III] luminosity above erg s-1, which is well above our sensitivity threshold). The median value of the halo mass for galaxies that are detected in the mocks is .
@@
In total, we draw 1000 mock galaxy samples (of each type) from the Uchuu simulations (stochastically accounting for the completeness function), which is the maximum number of fully independent mocks. As a validation of the mocks, we show that the distributions of [O III] luminosity and UV magnitude of the objects in the Uchuu mocks match relatively well the observed distributions in our sample (Fig. 7). This means that the [O III] LF and the [O III] - MUV relation in the mocks, the EIGER and C1F data-sets are consistent with each other. Generally, these distributions appear to be independent of the pointing direction choice (i.e. whether the mock is centering in a random position or towards a COLA1-like object).
We compare the redshift distribution of our sample of 141 [O III] emitters with the mocks described above. Due to resolution limitations of the Uchuu simulation, we compare the distribution of objects in our catalog with [O III] luminosity greater than erg s-1 with the average number of detected objects in the mocks. The peaks in the light gray histogram of Fig. 6 that are higher than the average number of detections in the mocks are over-densities. There is a peak of galaxies around the redshift of COLA1: we find 6 objects with in the sample, including COLA1 itself (4 objects if we limit the sample to erg s-1; )333, where number of detected objects in C1F in a given redshift interval, and the average number of detected objects in the mocks.. Interestingly, we find much more significant and stronger over-densities at other redshifts in the C1F (e.g., at , with , respectively). We analyse the over-density around COLA1 in detail in Section 5.
@@ 4 The properties of COLA1: all signs of a luminous Lyman Continuum leaker
In this section, we focus on the properties of COLA1 based on our JWST data, and previous Ly data that can now be better interpreted. We also investigate all other [O III] emitters in our dataset, with the intention to compare COLA1 with the galaxies in its environment, and quantify how special it is in comparison with these objects. Table 2 summarizes the main properties of COLA1.
@@ 4.1 Confirmation of the Ly systemic redshift
Cosmological radiation hydrodynamical reionization simulations predict a typical negligible transmission of the Ly blue peak for (e.g., Weinberger et al., 2018; Gronke et al., 2021). Hence, the presence of a strong blue peak at is very unlikely. The Ly profile of COLA1 is shown in Fig. 4. The main hypothesis described in Matthee et al. (2018) suggests that COLA1 is located in a highly ionised region (see also Hu et al. 2016), with a large enough size for the Ly photons to be redshifted out of resonance before reaching a relatively neutral IGM that would scatter them. Our measured systemic redshift using H and [O III] unambiguously confirm that the observed Ly profile is indeed a classic double-peaked Ly line, with the systemic redshift in between the two peaks (e.g., Verhamme et al., 2006; Gronke et al., 2015).
4.2 ISM conditions from optical emission lines
@@
The detection of the Balmer lines H and H permit to directly estimate the spectral attenuation due to dust. Assuming an intrinsic ratio of H/H (Miller, 1974), we measure for COLA1, whereas the value for the stack of the EIGER [O III] emitter sample at is , for comparison (Matthee et al., 2023a). Following a standard Cardelli et al. (1989) nebular attenuation curve (see also Reddy et al., 2020) leads to negligible dust attenuation in COLA1.
Combining the [O III]λ4364 fluxes of modules A and B we obtain a ¿ detection of [O III] erg s-1 cm-2. The detection of the [O III]λ4364 line in COLA1 allows for a direct measurement of the electron temperature, based on the [O III]λ4364/[O III]λ5008 ratio. For this we use the PyNeb Python package (Luridiana et al., 2015). For estimating the temperature, we assume an electron number density of cm-3 (e.g., Curti et al., 2023). We obtain K for COLA1. The measured for COLA1 is compatible with the typical values for galaxies at similar redshifts in the literature (e.g., Izotov et al., 2020; Matthee et al., 2023a; Katz et al., 2023; Nakajima et al., 2023; Hu et al., 2024).
We estimate the gas-phase metallicity of COLA1 from the [O III]/H ratio following Pilyugin et al. (2006). Since there is no direct measurement of any [O II] line, we assume a ratio of [O III]/[O II] , based on empirical results (e.g., Reddy et al., 2018; Katz et al., 2023) and following Matthee et al. (2023a). We obtain . The obtained metallicity is consistent with the average values observed in galaxies at similar redshifts (e.g., Nakajima et al., 2023). Under extreme ionization conditions, the [O III]/[O II] ratio can be much higher, however our metallicity estimation varies only slightly ( dex) if assumed ratios as high as (e.g., Topping et al., 2024).
4.3 UV properties
@@ 4.4 Star-formation rate and
The star-formation rate surface density is usually defined as where is the effective UV half-light radius of the galaxy (e.g., Naidu et al., 2020). In order to measure the effective radius of the C1F objects, we run SExtractor on the F200W image to obtain fwhm_world. Then, we stack 15 visually selected luminous stars, and fit a 2D Gaussian to the stack to estimate the average PSF full width at half maximum (FWHM). Finally we estimate the intrinsic FWHM as , valid for . We define the UV radius as the angular diameter distance corresponding to FWHM, according to our cosmology. The median value for the complete C1F sample is kpc. We find that COLA1 is unresolved in the F200W image, with FWHM arcsec, essentially indistinguishable from a stellar FWHM. Hence the above approximation is not valid for COLA1. This result only allows to set an upper bound to the UV size of COLA1 based on the stellar PSF, kpc. This UV size is smaller than the Ly size measured in Matthee et al. (2018), kpc, which is common in observed Ly emitters across a variety of redshifts (e.g., Leclercq et al., 2017).
@@ 4.5 Escape fraction of Ly and LyC photons
Several observational surveys and hydrodynamical simulations in the literature have investigated the correlation between the escape fraction of Ly and that of LyC photons (e.g., Verhamme et al., 2017). In Begley et al. (2024), a sample of 152 star forming galaxies with – is used to obtain a linear dependence between and . Using this relation we estimate, for COLA1, . In Maji et al. (2022), hydrodynamical and radiative transfer simulations are used to calibrate the same relation, suggesting ( if we correct for collisional Ly emission) for COLA1. Similarly, Kimm et al. (2022) fitted a power-law dependence from giant molecular cloud simulations, implying ( if we correct for collisional Ly emission) for COLA1.
@@
All the (LyC) estimations for COLA1 are summarized in Table 3. Apart from the calibration to the Ly escape fraction from Begley et al. (2024)666We note, however, that almost every object in the sample of Begley et al. (2024) has (Ly) , so our estimated LyC escape fraction based on this calibration is subject to a significant extrapolation., all estimates that are tied to COLA1s Ly emission indicate that it has a very high escape fraction of ionizing photons.

### v2

@@ 3.2 Redshift and line fluxes measurements
In order to measure the line fluxes and redshifts, we extract the 1D spectra of the candidates. We use the [O III]λ5008 line to trace the morphology of each object, as it is the most luminous emission line in our range. First, we select regions in the grism 2D EMLINE spectra containing the [O III]λ5008 line, and integrate them along the wavelength direction, for both modules A and B, whenever they are available. Then, we fit the spatial profile of the [O III]λ5008 emission with a single Gaussian for most cases, and double or triple Gaussian components for a few objects that present multiple clumps compatible with [O III] emission at the given redshift. Finally, we optimally extract the 2D spectra with the shape derived from the Gaussian fit to obtain the 1D spectrum of each candidate (Horne, 1986).
@@
Lastly, we measure the line fluxes. We fit individual Gaussian profiles to [O III]λ5008, [O III]λ4960 and H. We use narrow flat priors for the observed wavelengths of each line, based on the redshifts extracted from the [O III] doublets. In the particular case of COLA1, we also measure fluxes of the [O III]λ4364 and H, due to their exceptional S/N, as detailed in Sect. 3.3 below. Throughout this work, we use the average flux from the measurements of modules A and B (a single measurement is used for those sources where only one module is available).
3.3 Redshift calibration
@@ 3.4 SED modeling
We performed spectrophotometric SED fitting of all the [O III] emitters in our sample using the Prospector code (Johnson et al., 2021), including nebular emission and nebular continuum implemented via Cloudy (Ferland et al., 1998, 2013; Byler et al., 2017) and FSPS (Conroy et al., 2009, 2010; Conroy & Gunn, 2010) with the MIST stellar models (Choi et al., 2017). We follow the setup described in Naidu et al. (2022b), which in turn is adapted from Tacchella et al. (2022). We fit the F115W, F150W, F200W and F356W photometric fluxes, as well as H and [O III]λ5008,4960 line-fluxes simultaneously. The free parameters in the model include seven bins for the nonparametric star-formation history, with the first two being fixed to 0–5 Myr and 5–10 Myr (e.g., Tacchella et al. 2023), and the remaining logarithmically spaced out to . The other free parameters capture the total stellar mass formed, the stellar metallicity, the gas-phase metallicity, dust attenuation (a screen and an additional component for young stars), and the ionization parameter. We assume a Chabrier (2003) initial mass function, with a cutoff.
@@
Figure 5 shows the best-fit spectrum, the predicted photometry, and draws from the posterior that are compared to the observed data. The model predicts a best-fitting UV slope of , which is at odds with the UV photometry that shows a peculiar ”U” shape where the F150W flux lies below F115W and F200W (see Sect. 4.3). Keeping in mind that significant LyC leakage is at play, we fit a model with additional parameters that capture the escape fraction (Conroy & Kratter, 2012). In particular, a fraction of OB stars (uniform prior between 0 to ) is allowed to shine through the birth-cloud and dust-screen entirely unimpeded, which mimics “escape through holes;” in other words, an ionization-bounded nebula (e.g., Fig. 1 in Zackrisson et al. 2013). The predicted escape fraction from this fit is %. This high of is in agreement with the values obtained through different observational estimators in Sect. 4.5 below. However, due to the sparse data, we note the quality of our fit is effectively indistinguishable from a fit assuming and we defer investigation of this issue to future work, in particular when more photometric bands are available. This exploration highlights how COLA1 provides a unique test case, directly in the EoR, to hone estimators of LyC .
@@
COLA1 shows an average EW([O III]+H) when compared with [O III]-selected galaxies in the EIGER sample (Matthee et al., 2023). Using the grism fluxes of the [O III]+H lines and subtracting them from the F356W photometry to get a direct estimation of the continuum we get a value of EW0([O III]+H) Å, which is very consistent, with the Prospector prediction of EW0([O III]+H) Å. The fitted stellar mass is , but we note that these errors may be underestimated as they are partly bounded by the range in models that are included in Prospector. For example, if we would allow more recent and or exotic models, such as new implementations of binary star physics (Lecroq et al., 2024) or a top heavy initial mass function (e.g., Cameron et al., 2023a), it may well be possible that a somewhat lower stellar mass and star formation rate is required to reproduce COLA1s spectrum.
@@ 3.5 Galaxy mock catalogs in the COLA1 Field
In order to identify and quantify over-densities of the detected [O III] emitters in our data (shown in Fig. 6), we need to account for the selection function. As illustrated in Fig. 1, the field of view and sensitivity are both nontrivially dependent on wavelength. Moreover, another obvious factor in our selection function is that we targeted JWST to center on the bright galaxy COLA1. Because of these complications, we compare our observed distribution of galaxies to detailed mock observed samples of [O III] emitters that are derived by forward-modeling simulations through our selection function (i.e., the error cube that is illustrated in bottom panel of Fig. 1). This methodology has been developed by Mackenzie et al. (in prep.) in the context of understanding the environments of luminous quasars in the EIGER survey (see also Eilers et al. 2024), but is easily applicable to our survey.
As detailed in Mackenzie et al. (in prep.), large N-body simulations are combined with the UniverseMachine model (Behroozi et al., 2019) to obtain the SFR of each simulated halo. Then, by matching the rank order of the SFR and the [O III] luminosity (the property on which we select galaxies), one can derive an empirical description to obtain the [O III] luminosity of each simulated halo that is calibrated to match the observed [O III] luminosity function at . The observed [O III] luminosity function is an updated version of the one presented in Matthee et al. (2023), with minor changes. Since the clustering strength and the [O III] to UV luminosity relations of the simulated galaxies match the observed counterparts, this method seems to place the right galaxies in the right halos. The method is applied to the Uchuu (Ishiyama et al., 2021) dark matter only simulation. The Uchuu simulation is among the largest, (2 Gpc/)3, and therefore contains the rarest structures and large samples of halos hosting galaxies like COLA1, at the expense of resolution (i.e., it only resolves galaxies with [O III] luminosity above erg s-1, which is well above our sensitivity threshold). The median value of the halo mass for galaxies that are detected in the mocks is .
@@
In total, we draw 1000 mock galaxy samples (of each type) from the Uchuu simulations (stochastically accounting for the completeness function), which is the maximum number of fully independent mocks. As a validation of the mocks, we show that the distributions of [O III] luminosity and UV magnitude of the objects in the Uchuu mocks match relatively well the observed distributions in our sample (Fig. 7). This means that the [O III] LF and the [O III] - MUV relation in the mocks, the EIGER and C1F data-sets are consistent with each other. Generally, these distributions appear to be independent of the pointing direction choice (i.e., whether the mock is centering in a random position or towards a COLA1-like object).
We compare the redshift distribution of our sample of 141 [O III] emitters with the mocks described above. Due to resolution limitations of the Uchuu simulation, we compare the distribution of objects in our catalog with [O III] luminosity greater than erg s-1 with the average number of detected objects in the mocks. The peaks in the light gray histogram of Fig. 6 that are higher than the average number of detections in the mocks are over-densities. There is a peak of galaxies around the redshift of COLA1: we find 6 objects with in the sample, including COLA1 itself (4 objects if we limit the sample to erg s-1; )333, where number of detected objects in C1F in a given redshift interval, and the average number of detected objects in the mocks.. Interestingly, we find much more significant and stronger over-densities at other redshifts in the C1F (e.g., at , with , respectively). We analyze the over-density around COLA1 in detail in Section 5.
@@ 4 The properties of COLA1: all signs of a luminous Lyman Continuum leaker
In this section, we focus on the properties of COLA1 based on our JWST data, and previous Ly data that can now be better interpreted. We confirm the double-peak nature of the Ly line from the systemic redshift (Sect. 4.1), study the ISM conditions from the optical emission line measurements (Sect. 4.2) and the UV properties (Sect. 4.3); we estimate the star-formation rate and surface density (Sect. 4.4), and the escape fraction of Ly and LyC (Sect. 4.5). We also investigate all other [O III] emitters in our dataset, with the intention to compare COLA1 with the galaxies in its environment, and quantify how special it is in comparison with these objects. Table 2 summarizes the main properties of COLA1.
@@ 4.1 Confirmation of the Ly systemic redshift
Cosmological radiation hydrodynamical reionization simulations predict a typical negligible transmission of the Ly blue peak for (e.g., Weinberger et al., 2018; Gronke et al., 2021). Hence, the presence of a strong blue peak at is very unlikely. The Ly profile of COLA1 is shown in Fig. 4. The main hypothesis described in Matthee et al. (2018) suggests that COLA1 is located in a highly ionized region (see also Hu et al. 2016; Mason & Gronke 2020), with a large enough size for the Ly photons to be redshifted out of resonance before reaching a relatively neutral IGM that would scatter them. Our measured systemic redshift using H and [O III] unambiguously confirm that the observed Ly profile is indeed a classic double-peaked Ly line, with the systemic redshift in between the two peaks (e.g., Verhamme et al., 2006; Gronke et al., 2015).
4.2 Interstellar medium conditions from optical emission lines
@@
The detection of the Balmer lines H and H permit to directly estimate the spectral attenuation due to dust. Assuming an intrinsic ratio of H/H (Miller, 1974), we measure for COLA1 (H/H), whereas the value for the stack of the EIGER [O III] emitter sample at is , for comparison (Matthee et al., 2023). Following a standard Cardelli et al. (1989) nebular attenuation curve (see also Reddy et al., 2020) leads to negligible dust attenuation in COLA1.
Combining the [O III]λ4364 fluxes of modules A and B we obtain a ¿ detection of [O III] erg s-1 cm-2. The detection of the [O III]λ4364 line in COLA1 allows for a direct measurement of the electron temperature, based on the [O III]λ4364/[O III]λ5008 ratio. For this we use the PyNeb Python package (Luridiana et al., 2015). For estimating the temperature, we assume an electron number density of cm-3 (e.g., Curti et al., 2023). We obtain K for COLA1. The measured for COLA1 is compatible with the typical values for galaxies at similar redshifts in the literature (e.g., Izotov et al., 2020; Matthee et al., 2023; Katz et al., 2023; Nakajima et al., 2023; Hu et al., 2024).
We estimate the gas-phase metallicity of COLA1 from the [O III]/H ratio following Pilyugin et al. (2006). Since there is no direct measurement of any [O II] line, we assume a ratio of [O III]/[O II] , based on empirical results (e.g., Reddy et al., 2018; Katz et al., 2023) and following Matthee et al. (2023). We obtain . The obtained metallicity is consistent with the average values observed in galaxies at similar redshifts (e.g., Nakajima et al., 2023). Under extreme ionization conditions, the [O III]/[O II] ratio can be much higher; however, our metallicity estimation varies only slightly ( dex) if assumed ratios as high as (e.g., Topping et al., 2024b).
4.3 Ultraviolet properties
@@ 4.4 Star-formation rate and
The star-formation rate surface density is usually defined as where is the effective UV half-light radius of the galaxy (e.g., Naidu et al., 2020). In order to measure the effective radius of the C1F objects, we run SExtractor on the F200W image to obtain fwhm_world. Then, we stack 15 visually selected luminous stars, and fit a 2D Gaussian to the stack to estimate the average PSF full width at half maximum (FWHM). Finally we estimate the intrinsic FWHM as , valid for . We compute the UV radius as the physical distance corresponding to FWHM666Here we assume that the light profile of COLA1 is described by a Gaussian (Sérsic profile with ). However, the measurement of the effective radius could vary under the assumption of different Sérsic indices.. The median value for the complete C1F sample is kpc. We find that COLA1 is unresolved in the F200W image, with FWHM arcsec, essentially indistinguishable from a stellar FWHM. Hence the above approximation is not valid for COLA1. This result only allows to set an upper bound to the UV size of COLA1 based on the stellar PSF, kpc. This UV size is smaller than the Ly size measured in Matthee et al. (2018), kpc, which is common in observed Ly emitters across a variety of redshifts (e.g., Leclercq et al., 2017).
@@ 4.5 Escape fraction of Ly and LyC photons
Several observational surveys and hydrodynamical simulations in the literature have investigated the correlation between the escape fraction of Ly and that of LyC photons (e.g., Verhamme et al., 2017). In Begley et al. (2024), a sample of 152 star forming galaxies with – is used to obtain a linear dependence between and inferred from H measurements assuming a Ly/H ratio. Using this relation we estimate, for COLA1, . In Maji et al. (2022), hydrodynamical and radiative transfer simulations are used to calibrate the same relation, suggesting ( if we correct for collisional Ly emission) for COLA1. Similarly, Kimm et al. (2022) fitted a power-law dependence from giant molecular cloud simulations, implying ( if we correct for collisional Ly emission) for COLA1.
@@
The Ly central escape fraction () is introduced in Naidu et al. (2022a) as a diagnostic of LyC leakage, defined as the fraction of Ly flux emitted within km s-1 from the systemic Ly velocity. The Ly central fraction traces the presence of low opacity channels that allow the direct escape of Ly. From the Ly spectrum of COLA1 (see Fig. 4) we obtain %. Recent radiative-transfer simulations have found that sources with are more likely to be LyC leakers (Choustikov et al., 2024). Direct comparison with the ”High Escape” stack in Naidu et al. (2022a) suggests .
All the (LyC) estimations for COLA1 are summarized in Table 3. Apart from the calibration to the Ly escape fraction from Begley et al. (2024)777We note, however, that almost every object in the sample of Begley et al. (2024) has (Ly) , so our estimated LyC escape fraction based on this calibration is subject to a significant extrapolation., all estimates that are tied to COLA1s Ly emission indicate that it has a very high escape fraction of ionizing photons.